# Mon CHÉRI: Mitigating Uninitialized Memory Access with Conditional Capabilities

### v2

@@ A.1 Example of Avoided Data Hazard
Listing 2: Example of instruction sequence causing data hazard as described in § 4. At ➀, the csetwbrbound instructions sets the Write-before-Read bound for the capability in register ca0, turning it into a conditional capability. At ➁, the capability store word (csw) instructions performs a store operation on ca0, whereupon the operation bound of the capability in ca0 is increased at Stage-2 of the pipeline. At the same time, the capability load word (clw) instruction at ➂ has already
@@ A.2 Changes to Capability Encoding
[…]
$E&=0\\ T\left[2:0\right]&=T_{E}{}\\ B\left[2:0\right]&=B_{E}{}\\ \leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}O\left[2:0\right]&\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}=\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}O_{E}\left[4:2\right]\\ L_{carry\_out}&=\begin{cases}\begin{aligned} 1,\quad&\textrm{if }T\left[11:0\right]<B\left[11:0\right]\\ 0,\quad&\textrm{otherwise}\end{aligned}\end{cases}\\ L_{msb}&=0\\ \end{aligned}$ $E&=\left\{T_{E},B_{E}\right\}\\ T\left[2:0\right]&=0\\ B\left[2:0\right]&=0\\ \leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}O\left[2:0\right]&\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}=\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}O_{E}\left[E+2:E\right]\\ L_{carry\_out}&=\begin{cases}\begin{aligned} 1,\quad&\textrm{if }T\left[11:3\right]<B\left[13:3\right]\\ 0,\quad&\textrm{otherwise}\end{aligned}\end{cases}\\ L_{msb}&=1\\ \end{aligned}$
Reconstituting the top two bits of $T$:
$\begin{split}T\left[13:12\right]&=B\left[13:12\right]+L_{carry\_out}+L_{msb}\\ \end{split}$
@@
address, $a=$ $a_{top}=a\left[\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}47:E+14\right]$ $a_{mid}=a\left[E+13:E\right]$ $a_{low}=a\left[E-1:0\right]$
top, $t=$ $a_{top}+c_{t}$ $T\left[13:0\right]$ $0^{\prime}E$
@@
$\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}O_{3}$ $=\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}O\left[13:11\right]$ (4)
@@
false false 00
false true $+1$
true false $-1$
true true 00
@@
false false 00
false true $+1$
true false $-1$
true true 00
@@
false false 00
false true $+1$
true false $-1$
true true 00

### v3

@@ A.1 Example of Avoided Data Hazard
1/* ... */
2  csetbounds ca0, ca0, 4
3csetwbrbound ca0, ca0, zero
4  li a1, 10
5csw a1, 0(ca0)
6clw a0, 0(ca0)
7  clc cra, 16(csp)
8/* ... */
Listing 2: Example of instruction sequence causing data hazard as described in §˜4. At ➀, the csetwbrbound instructions sets the Write-before-Read bound for the capability in register ca0, turning it into a conditional capability. At ➁, the capability store word (csw) instructions performs a store operation on ca0, whereupon the operation bound of the capability in ca0 is increased at Stage-2 of the pipeline. At the same time, the capability load word (clw) instruction at ➂ has already
@@ A.2 Changes to Capability Encoding
023131416172526274445464748596063
$p_{op}$
$p_{hw}$
$f$
otype
$I_{E}$
$T\left[11:3\right]$
$T_{E}$
$B\left[13:3\right]$
$B_{E}$
$O\left[13:3\right]$
$O_{E}$
$a$
Cursor
@@
address, $a=$ $a_{top}=a\left[\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}47:E+14\right]$ $a_{mid}=a\left[E+13:E\right]$ $a_{low}=a\left[E-1:0\right]$
top, $t=$ $a_{top}+c_{t}$ $T\left[13:0\right]$ă $0^{\prime}E$
@@
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}O_{3}$ $=\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}O\left[13:11\right]$ (4)
@@
false false 0
false true +1+1
true false 1-1
true true 0
@@
false false 0
false true +1+1
true false 1-1
true true 0
@@
false false 0
false true +1+1
true false 1-1
true true 0